OpenStudy (anonymous):

need help on this problem 5x-y=-9 x-y=-1 need help please really quick ill give a medal

OpenStudy (anonymous):

y = 5x + 9 x - 5x + 9 = -1 -4x + 9 = -1 -4x = -10 x = 5/2

OpenStudy (anonymous):

(1,2) (2,1) (-1,-2) (-2,-1) these are my choices

OpenStudy (anonymous):

(1,2) works only for the second one.

OpenStudy (anonymous):

It's the last one (-2,-1)

OpenStudy (anonymous):

so which of my choices would fit work for this problem